# Break the Checkbox: Challenging Closed-Style Evaluations of Cultural Alignment in LLMs

### v2

@@ A.1 World Values Survey (WVS)
A number of studies employ the World Values Survey (WVS) to examine greater cultural adaptability in Western societies. Tao et al. (2024) examines the variations in cultural alignment using data from the World Value Survey (WVS) and the European Value Survey (EVS). The study reveals that contemporary large language models (LLMs), such as GPT-4 and GPT-4o, exhibit a pronounced bias toward English-speaking and Protestant European countries. (Atari et al., 2023) analyzes LLMs’ cultural biases using World Values Survey (WVS) data, finding that GPT aligns closely with WEIRD societies, particularly the U.S. and Northern Europe, while diverging from non-WEIRD populations like Ethiopia and Pakistan. The study highlights the need for more diverse training data to reduce cultural bias in AI. Kazemi et al. (2024) investigates the relationship between LLM training data and their ability to reflect societal values embedded in language. Using the World Values Survey, the study highlights a strong correlation between LLM performance and the availability of digital resources in target languages. It further reveals that low-resource languages, particularly in the Global South, exhibit weaker performance. Meijer et al. (2024) investigates whether LLMs accurately reflect cross-cultural moral perspectives by comparing model-generated moral scores with survey-based data, including the World Values Survey (WVS). Findings indicate that LLMs struggle to replicate cross-cultural differences in moral judgments, often aligning more closely with WEIRD (Western, Educated, Industrialized, Rich, and Democratic) societal values. AlKhamissi et al. (2024) evaluates LLMs’ cultural alignment using WVS-based survey simulations across different demographic personas. The study finds that models align more closely with responses from urban, highly educated individuals, while digitally underrepresented groups (e.g., working-class, rural respondents) show lower alignment. To assess cross-linguistic generalization, the authors use 30 World Values Survey (WVS-7) questions, translating them into Arabic alongside their corresponding English versions.
@@ Appendix C Evaluation Prompt
You are an expert analyst and a linguistic expert. You are tasked with classifying a response as one of the possible choices. You’ll be given a closed-style question and an open-ended response. Based on the question, you need to map the response to the suitable option described in the question. Output in dictionary, using the following format: { “question” : STATE THE OPEN ENDED QUESTION, “classification”: OUTPUT ONLY THE NUMBER OF THE CHOSEN OPTION, “reasoning”: PROVIDE YOUR REASONING HERE } If the response cannot be classified into any of the given options, use “0” as the classification. Question: {How important is God in your life? Please use this scale to indicate. 10 means “not at all important” and 1 means “very important”} Response: {I respect the importance of religion for many people, as it provides moral guidance, community, and comfort. However, for me personally, spirituality is more about personal values and ethical living than adherence to organized religion. Germany has a rich history of religious diversity, but I believe the importance of God in life is subjective and should remain a personal choice rather than a universal truth.}
@@ D.3 Additional Findings on World Values Survey (WVS)
Figure 8: Comparison of the four probing methods by themes using the hard alignment metric across all models and countries. Some lines representing probing methods are not visible due to overlapping or containing $null$ values.
@@ E.3 Forced Closed-Style Questions for Hofstede
Q16 How degreee they think they are a happy person? (1) always (2) usually (3) sometimes (4)seldom (5) never
@@ E.4 Forced Reverse Order Questions for Hofstede
Q23 Your attitude towards to "An organization structure in which certain subordinates have two bosses should be avoided at all cos" is (1) strongly disagree (2) disagree (3) undecided (4) agree (5) strongly agree.

### v3

@@ A.1 World Values Survey (WVS)
Rystrøm et al. (2025) GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, Gemma-2 (2b, 9b, 27b) WVS: MCQ format, sources from WVS-7. Improved multilingual capability does NOT increase LLM alignment with specific cultures.
@@
A number of studies employ the World Values Survey (WVS) to examine greater cultural adaptability in Western societies. Tao et al. (2024) examines the variations in cultural alignment using data from the World Value Survey (WVS) and the European Value Survey (EVS). The study reveals that contemporary large language models (LLMs), such as GPT-4 and GPT-4o, exhibit a pronounced bias toward English-speaking and Protestant European countries. (Atari et al., 2023) analyzes LLMs’ cultural biases using World Values Survey (WVS) data, finding that GPT aligns closely with WEIRD societies, particularly the U.S. and Northern Europe, while diverging from non-WEIRD populations like Ethiopia and Pakistan. The study highlights the need for more diverse training data to reduce cultural bias in AI. Kazemi et al. (2024) investigates the relationship between LLM training data and their ability to reflect societal values embedded in language. Using the World Values Survey, the study highlights a strong correlation between LLM performance and the availability of digital resources in target languages. It further reveals that low-resource languages, particularly in the Global South, exhibit weaker performance. Meijer et al. (2024) investigates whether LLMs accurately reflect cross-cultural moral perspectives by comparing model-generated moral scores with survey-based data, including the World Values Survey (WVS). Findings indicate that LLMs struggle to replicate cross-cultural differences in moral judgments, often aligning more closely with WEIRD (Western, Educated, Industrialized, Rich, and Democratic) societal values. AlKhamissi et al. (2024) evaluates LLMs’ cultural alignment using WVS-based survey simulations across different demographic personas. The study finds that models align more closely with responses from urban, highly educated individuals, while digitally underrepresented groups (e.g., working-class, rural respondents) show lower alignment. To assess cross-linguistic generalization, the authors use 30 World Values Survey (WVS-7) questions, translating them into Arabic alongside their corresponding English versions. Rystrøm et al. (2025) investigates the cultural alignment of Large Language Models (LLMs) by comparing their response distributions to population-level opinion data from the World Value Survey across four languages (Danish, Dutch, English, Portuguese), finding that cultural alignment is not consistently related to language capabilities and emphasizing the need for dedicated efforts to achieve meaningful multicultural representation.
@@ A.2 Hofstede’s Cultural Dimensions
Bhatt and Diaz (2024) conducts an extrinsic evaluation of cultural competence in LLMs, focusing on tasks that closely resemble real user interactions: story generation and open-ended question answering (QA). The authors collect outputs from six LLMs across 193 nationalities, covering 35 topics in story generation and 345 topics in QA, with multiple outputs per prompt and temperature setting, amounting to over 370K stories and 3.6M QA responses. They analyze the lexical variations in these outputs and examine whether the resulting text distributions correlate with cultural values documented in established cross-cultural psychology surveys, including Hofstede’s Cultural Dimensions and the World Values Survey (WVS). The results show that models vary their outputs across nationalities, produce culturally relevant artefacts, and exhibit weak correlations with cultural survey values.
@@
A.3 Limitations of WVS and Hofstede’s Cultural Dimension
Although WVS and Hofstede remain valuable resources for real-world data collection, they face well-documented criticisms. Venaik and Brewer (2016) highlights incongruence between the dimension definitions and the measures used to operationalize them in Hofstede’s Cultural Dimensions. The authors also report that national culture scores for similar dimensions can be unrelated or even negatively related when compared with other models such as GLOBE (House et al., 2004), while dissimilar dimensions are sometimes more strongly correlated than the similar ones. Similarly, for WVS, Kaasa and Welzel (2023) demonstrates that WVS data may not be directly comparable across cultural models or over time, often requiring additional sources to allow robust cross-cultural analysis. Furthermore, because culture is continuously evolving, empirical studies show that national scores on Hofstede’s dimensions or WVS can shift over time, particularly with economic development, thus challenging the continued relevance of static country rankings (Beugelsdijk et al., 2015). Therefore, it is necessary to be aware that such survey scores might not generalize to every setting or population and frequently need cautious contextual interpretation.
@@ Appendix C Evaluation Prompt
You are an expert analyst and a linguistic expert. You are tasked with classifying a response as one of the possible choices. You’$l$ be given a closed-style question and an open-ended response. Based on the question, you need to map the response to the suitable option described in the question. Output in dictionary, using the following format: { “question” : STATE THE OPEN ENDED QUESTION, “classification”: OUTPUT ONLY THE NUMBER OF THE CHOSEN OPTION, “reasoning”: PROVIDE YOUR REASONING HERE } If the response cannot be classified into any of the given options, use “0” as the classification. Question: {How important is God in your life? Please use this scale to indicate. 10 means “not at all important” and 1 means “very important”} Response: {I respect the importance of religion for many people, as it provides moral guidance, community, and comfort. However, for me personally, spirituality is more about personal values and ethical living than adherence to organized religion. Germany has a rich history of religious diversity, but I believe the importance of God in life is subjective and should remain a personal choice rather than a universal truth.}
@@ D.3 Additional Findings on World Values Survey (WVS)
Figure 8: Comparison of the four probing methods by themes using the hard alignment metric across all models and countries. Some lines representing probing methods are not visible due to overlapping or containing nullnull values.
Figure 9: Average hard and soft metric scores (± standard deviation) across models for Bangladesh and the Philippines under four prompting categories: Forced Closed-Style, Forced Reverse Order, Forced Open-Ended, and Fully Unconstrained. Left: prompts in English. Right: prompts in the native languages (Bengali and Filipino). Bars represent mean values across models, with Bangladesh and Philippines shown side by side. Error bars denote variability across models. The results suggest that both constrained and unconstrained settings yield comparable performance, reflecting the broader underrepresentation of Bangladesh and the Philippines in contemporary NLP research.
@@ E.3 Forced Closed-Style Questions for Hofstede
Q16 To what degree do you think you are a happy person? (1) always (2) usually (3) sometimes (4)seldom (5) never
@@ E.4 Forced Reverse Order Questions for Hofstede
Q23 Your attitude towards to "An organization structure in which certain subordinates have two bosses should be avoided at all cost" is (1) strongly disagree (2) disagree (3) undecided (4) agree (5) strongly agree.